OpenStudy (jordans_openstudy):

Write the equation of a line parallel to the line 2x+3y=5 and passes through the point (9,-3)

OpenStudy (luigi0210):

First get your equation in slope-intercept form: \(y=mx+b\) and m will be your slope. An equation that is parallel to it will have the same slope. Then just use the point-slope formula: \(\large y-y_1=m(x-x_1)\)

OpenStudy (jordans_openstudy):

not really sure :/

OpenStudy (texaschic101):

do you know how to put the equation 2x + 3y = 5 in y = mx + b form ?

OpenStudy (jordans_openstudy):

Nope :l i am so bad at math lol

OpenStudy (texaschic101):

2x + 3y = 5.......we are basically solving for y subtract 2x from both sides 2x - 2x + 3y = -2x + 5 combine like terms 3y = -2x + 5 divide both sides by 3 (3/3)y = (-2/3)x + 5/3 y = -2/3x + 3 read ^^ and see if you have any questions ? Then we will continue....

OpenStudy (texaschic101):

well.....questions ? or shall we continue ?

OpenStudy (jordans_openstudy):

hmm i think I'm starting to get it

OpenStudy (texaschic101):

oops....I meant y = -2/3x + 5/3...I forgot the 5

OpenStudy (jordans_openstudy):

no worries haha, yea thats a option it shows

OpenStudy (texaschic101):

so our equation in y = mx + b form is : y = -2/3x + 5/3 In y = mx + b form, the slope is the number in the m position. So our slope is -2/3 Parallel lines WILL HAVE same slope. still with me ?

OpenStudy (jordans_openstudy):

ohhh okay

OpenStudy (texaschic101):

now we take y = mx + b slope (m) = -2/3 your points (9,-3)....x = 9 and y = -3 so basically the only thing we need to find is b, the y intercept y = mx + b -3 = -2/3(9) + b -3 = -6 + b -3 + 6 = b 3 = b so our parallel equation is : y = -2/3x + 3

OpenStudy (texaschic101):

we all good ? I tried to explain as best I could, but if you have any questions, I can try to explain better.

OpenStudy (jordans_openstudy):

Were all good thanks i appreciate it :)